A Complete Bibliography of Publications in Revue d’Histoire Mathématiques = Journal

for the History of Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 April 2021 Version 1.03

Title word cross-reference

XVI^{e} [198].

10106 [140]. 15 [154]. 17th [188, 13]. 1920s [173, 157]. 1930s [21]. 19th [219, 189].

2000g [37]. 2001a [65].

4 [37, 65].

5e [47]. 5th [47].

Âge [123].

Abbaco [97]. Académie [138, 99]. Academy [138, 99]. according [187, 150]. activité [116]. activity [116]. adaptation [204]. added [96]. Addendum [154]. advances [198]. Ages [123]. agitateur [56]. agitator [56]. aircraft [194]. Alî [174]. Alfred [213]. algebra [48, 92, 105, 30, 117, 107, 62, 218]. Algebraic [17, 134, 164, 200, 218, 203]. algebraischen [218]. Algebras [204]. Algèbre [62, 105, 30, 107]. algèbres [204]. algébrique [134]. algébriques [200]. algébristes [198]. algebrists [198]. alone [7]. always [151]. analyse [134, 5]. analysis [134, 205, 5, 108, 117]. Analytical [179]. analytique [179]. ancient [59]. Anglais [72]. angle [72]. annales [220]. années [21, 110]. Annuities [78]. Anti [194]. Anti-aircraft [194]. antiquité [79]. antiquity [79, 37, 31]. apercu [8, 199]. Apollonius [35]. applied [14]. appliquées [14]. approchée [127]. approximate [127]. approximation [127]. Arabic [111]. Archimedean [122]. architectural [145]. area [46, 96]. arguments [188]. Arismetica [192]. Aristotelian [77]. aristotélicienne [77]. arithmetic [174, 106, 140, 107]. Arithmetica [203]. Arithmetician [29]. arithméticien [29]. arithmetics [192]. arithmétique [107]. arithmétiques [174, 192, 106]. art [46, 212]. Article [91, 93, 91]. Ascalon [35]. Asset [214]. astronomie [181, 171]. astronomy [205, 83, 181, 171]. attraction [11]. attribués [174]. attributed [174, 180]. authoritarian [99]. Authorship [216]. automne [217]. autoritaire [99]. Autour [10, 87, 3, 138, 57]. autre [69]. autres [174, 2]. aux [14, 139, 21, 204, 189]. avancées [198]. avant [2]. axiom [69]. axiome [69]. axiomes [52]. axioms [52].

B [65]. Babylonian [96, 54]. babylonienne [54]. Bachelier [214, 176]. Bailly [83]. Baire [2]. Barcelona [195]. barcelonais [195]. based [7, 110]. be [112]. before [2, 117]. Belhoste [65]. Berlin [63]. between [156, 121, 155, 217, 92, 17, 46, 30, 72, 118, 52, 117]. Betwixt [83]. Bézout [134]. Bibliographic [191, 190]. bibliographiques [190]. Biblioteca [140]. Bibliothèque [105]. Bidwell [157]. Bienaymé [20]. bigger [151]. Birkhoff [42]. birth [149]. board [190, 191]. Bohuslav [118]. Book [47, 36, 58]. Books [189]. borderland [46]. Borel [127, 41, 173, 2, 103]. Bos [200]. boundary [11]. boundary-value [11]. Bourbaki [110]. Bourdon [204]. boycott [157]. Branch [219]. Britain [205]. British [184]. Bürgi [172].

cadastres [76]. cadres [53]. cahiers [210]. Calcul [146, 20, 168, 88, 54, 49]. calculate [133]. calculation [20, 146, 88, 54]. calculer [133]. Calculus [214, 146]. called [42]. camels [174]. Camille [162, 121, 163]. Campanus [123, 36]. Can [51]. candidat [135]. candidate [135]. caractéristiques [218]. Carnot [28]. Cartan [144, 210]. Cartesian [62]. cartésiennes [62]. cartographie [210]. cartography [210]. cas [138, 144]. Case [220, 138, 144, 43, 175]. categories [110, 110]. Cauchy [24, 146]. céleste [193]. celestial [193]. center [88]. centesimal [168, 168]. centre [88]. centuries [184]. Century [189, 134, 8, 53, 67, 219, 105, 188, 13, 6, 198, 197, 86]. cercle [99]. certaines [116]. chaînon [156]. chaleur [179]. chameaux [174]. chapter [17]. characteristics [218]. Characterizing [211]. Charles [221, 161]. Chavanne [162]. Chess [71]. chez [72, 4, 3]. chiite [174]. Chilias [215]. choice [69]. choix [69]. Cholesky [104]. Chuquet [107]. cinquante [110]. circle [99, 188]. Circulation [216, 174, 203]. Civita [109]. Classical [46]. Clebsch [213]. cloth [71]. combinations [98]. come [117]. comité [190]. Comment [79, 65]. Commentaries [37]. Commentatio [185]. commercial [106, 140, 107]. commerciale [107]. community [34]. compagnie [106]. company [106]. comparaison [212]. comparative [91]. comparison [212]. compendium [47]. Complement [200, 200]. complex [5]. complexe [5]. compromis [204]. Compromise [204]. computing [194, 168]. concept [79, 63, 15]. conception [87]. Conceptions [48]. Conics [35]. Coniques [35]. conjecture [200]. connaissance [127]. connection [205]. constructive [9]. contact [72]. context [192, 107]. contexte [192, 107]. continue [123]. continuous [123]. contrôle [196]. Contribution [6, 10, 6]. contributions [122, 109]. control [196]. controverse [121, 179]. controversy [121, 179, 52]. convergence [184]. convex [93]. convexes [93]. cordes [179]. correspondance [199]. correspondence [199]. couldn’t [73]. courbes [200, 213]. Cournot [4]. cours [63, 110]. course [63]. covariant [15]. creation [164, 208, 208]. crise [2]. crisis [2]. criticized [20]. critiquée [20]. crossroads [197]. Cubic [213]. cubiques [213]. culture [97]. curieuses [71]. Curious [71]. Curves [213, 219, 200]. cusp [64].

d’Alembert [139]. dans [32, 192, 143, 145, 139, 21, 179, 106, 171, 88, 77, 212, 215, 36, 123, 193, 107, 116, 218]. d’Apollonios [35]. d’après [7]. d’architecture [145]. d’arismetica [192]. d’Ascalon [35]. Davenport [143]. day [194]. decision [99, 99]. decoration [46]. d’écriture [93]. define [79]. définir [79]. definite [188]. definition [18, 123, 18, 123]. definitions [209]. degré [198]. degree [68, 198]. Dehn [122]. del [128]. d’Élie [144, 210]. d’Émile [127]. demonstrations [19, 19]. d’enseignement [189]. derivation [18, 18]. dérivées [139]. dernier [8, 86]. Descartes [170, 200, 112]. design [87]. determinant [175]. d’Euclide [77, 58, 123, 47]. Deuteronomic [31, 79, 37]. deutéronomiques [79]. Deux [76, 20]. development [10]. développement [10]. d’Hermite [10]. d’histoire [190, 191, 154]. diagrammes [171]. diagrams [171]. Dickson [40]. didactic [87]. didactique [87]. Differential [164, 139, 108, 23]. differentiation [15]. différentielles [23]. difficulté [123]. difficulty [123]. d’Irenée [20]. discipline [143]. discovery [172]. discussion [8]. dit [28, 112]. diter [176]. dix [174]. dix-sept [174]. do [117]. documents [214, 221]. Doeblin [118]. Doetsch [34]. Doing [26]. Domestic [204]. domestiques [204]. d’ordre [18]. Dugac [44]. d’un [10, 51, 57, 93, 3]. D’une [5, 208, 18, 210, 13]. d’Ur [76]. durant [86, 208]. during [86, 208]. Duvillard [78]. dynamics [12]. dynamique [12]. early [172, 157, 105]. Échiquiers [71]. école [9, 135]. écoulement [91]. écrite [54]. écrits [145, 88]. écriture [91]. édité [35]. edited [35]. Éditer [176]. edition [123, 123]. editions [37]. Editorial [167, 178, 183, 202, 207, 75, 81, 85, 90, 95, 102, 115, 120, 125, 153, 131, 137, 148, 142, 158, 190, 191, 166, 177, 182, 201, 206, 124, 22, 27, 33, 39, 45, 50, 55, 61, 66, 70, 74, 80, 84, 89, 94, 101, 114, 119, 152, 130, 136, 147, 141, 159]. Editors [132]. Edmund [157]. Édouard [29]. education [32, 53]. Edwin [157]. efforts [164]. eighteenth [184]. elaboration [18, 18]. élémentaire [208]. elementary [208]. Elements [111, 36, 58, 123, 47, 36, 58, 123, 47]. Élie [144, 210]. elle [110]. elliptic [62, 10]. Elliptical [213]. elliptique [213]. elliptiques [62, 10]. emergence [173, 13, 13]. Émile [127]. eminence [172]. End [189, 8]. engineer [196]. engineering [196]. English [72]. Enlightened [134, 99, 181]. enlightenment [83]. enquêteur [56]. enseignement [32, 53, 208]. ensembles [2]. entier [18]. entre [156, 121, 155, 217, 92, 30, 72, 52]. enumerators [217]. epistemology [127]. eponymy [175]. époque [76]. equal [209]. equation [68]. equations [164, 139, 24, 108, 212, 198, 23, 139, 24, 212, 198, 23, 218]. equimultiples [123, 123]. era [208]. ère [208]. esopotamia [168]. espace [195, 204]. espaces [16]. Espagnes [106]. Esquisse [210]. essay [7]. essor [14]. est [151]. est-il [151]. États [14, 204]. États-Unis [14, 204]. Étienne [134]. Euclid [47, 111, 77, 58, 123]. Europe [212]. European [198]. européens [198]. Eutocius [35]. Évariste [163, 135, 160, 165]. everything [151]. evolution [11]. examples [143]. exchanges [118]. exciting [111]. exemples [143]. existence [11]. exploits [174]. Exploration [93, 93]. expressability [58]. exprimabilité [58].

F. [57]. Fīrūz [111]. Factor [149]. factoring [57]. factorisation [57]. Faire [26]. fall [217]. fascicule [74]. Feldbau [138]. Fermat [126]. Fibonacci [97]. ficelle [116]. fictions [218]. fields [161]. fifteenth [105]. fifties [110]. figures [35]. fin [8, 189]. Finance [214]. first [19, 16, 199, 217, 106, 57]. Five [216]. fixe [3]. fixed [3]. Flow [91]. Fluid [91, 67]. fluides [91]. fonction [170]. fonctions [62, 10, 16, 5]. fonde [110]. fonde-t-elle [110]. fondements [150]. forbid [99]. forgotten [17]. form [7]. former [7]. formula [19]. Formulario [211]. formule [19]. foundations [150]. Fourier [19, 179]. frameworks [53]. francais [217, 107]. francaises [204]. France [8, 86, 2, 87, 105]. Francesco [47]. Francisco [221]. Francois [87]. Fréchet [56]. French [154, 69, 174, 134, 19, 190, 164, 56, 138, 127, 62, 10, 32, 79, 195, 78, 63, 156, 121, 104, 20, 41, 155, 128, 173, 168, 8, 86, 28, 144, 192, 53, 208, 205, 14, 29, 71, 35, 16, 18, 186, 199, 217, 135, 7, 92, 113, 143, 145, 2, 91, 139, 210, 21, 170, 179, 200, 99, 110, 51, 106, 87, 105, 30, 213, 171, 5, 72, 24, 146, 88, 9, 112, 4, 103, 77, 13, 57, 212, 6, 52, 26, 196, 204, 54, 133, 76, 93, 189, 187, 198, 215, 36, 58]. French [123, 193, 3, 107, 47, 23, 49, 82, 116, 176, 218, 12, 150, 151, 181, 165]. Friedrich [150]. Function [170]. functional [108]. functions [62, 10, 16, 5, 139].

G [128, 108]. Gaitapañcaviśī [180]. Galilean [12]. galiléenne [12]. Galois [164, 156, 163, 135, 160, 161, 165]. games [116]. Gaston [217]. Gauss [156]. general [57]. générale [57]. généralité [28, 93]. generality [28, 93]. Genèse [16, 200, 103]. genesis [15, 200, 103, 16]. geodesic [6, 93]. géodésiques [6, 93]. Geometria [216, 129]. Geometrical [46]. Géométrie [71, 62, 28, 208, 92, 143, 145, 21, 88, 77, 52, 189, 150, 113, 170]. geometries [122]. géométrisation [186]. geometrization [186]. Geometry [189, 62, 28, 208, 92, 143, 145, 17, 21, 209, 88, 77, 96, 52, 150, 71, 143, 113, 170]. Geppert [217]. German [157]. Germany [34, 197]. Gleichungen [218]. glimpse [199]. global [144]. Gödel [128]. Going [144]. Gomes [221]. Gramegna [108]. grand [151]. graphic [86]. graphical [82]. graphing [49]. graphique [86, 49, 82]. gravité [88]. gravity [88]. great [205]. grecque [79]. Greek [79, 46, 25, 197]. ’Grothendieck [110, 110]. guerre [14]. guns [194]. Gustav [34].

H. [200]. Hamilton [3]. handle [145]. Harald [217]. Harriot [98]. Hasse [217]. Heat [179]. Hecke [199]. Hecke/Weyl [199]. Helmut [217]. Henri [173]. Hermite [10, 221, 161]. Hilbert [122]. Histoire [37, 65, 32, 143, 6, 49]. historical [40, 200, 30]. historiography [83, 25]. historique [200, 30]. History [222, 223, 224, 32, 37, 172, 143, 17, 6, 31, 60, 83, 65, 197, 49, 40]. Hobbes [72]. homologie [21]. homology [21]. Hostinský [118]. house [176]. Huygens [12]. hydrodynamics [109].

Iberian [192, 140, 203]. ibériques [192]. ideal [209]. III [76]. imam [174, 174]. impact [122, 14, 59]. implications [40]. impossibility [188]. imprimées [106]. Inaugural [170]. inaugurale [170]. inauguration [179, 200]. incidence [92]. indefinite [188]. indeterminate [68]. Indian [83, 181]. indicateurs [190]. indicators [190, 191]. indienne [181]. inédit [10, 63]. inédits [20]. ingénierie [196]. ingénieur [196]. inquiry [46]. inspired [37]. instability [67]. Institut [173]. Institutional [53]. institutionnels [53]. instrumentation [29]. instruments [145]. integral [19, 146, 146]. intégrale [19]. integration [23, 82, 23, 82]. Interdire [99]. intermédiaires [189]. Intermediate [189]. internal [78]. interne [78]. interpretations [38]. intertextuality [100]. introduction [179, 200]. Inventive [38]. invert [97]. investigator [56]. Ireland [205]. Irenée [20]. Irrationalité [187]. irrationality [187, 187]. Italian [122, 145, 198]. italiens [145, 198].

J [43]. J. [43]. Jacques [138]. Japanese [208, 60]. japonaise [208]. Jean [83]. Jean-Sylvain [83]. Jesuit [83]. jeux [116]. John [88]. Joint [132]. Jordan [162, 121, 163]. Journals [132, 220, 189, 184]. Jules [20]. Julia [217]. Junius [82].

Kaapayādi [169]. Karl [194]. Kepler [215]. Kerala [169]. Knowing [145]. Knowledge [211, 216, 127, 203]. Kronecker [63, 121, 218].

Lacroix [48, 87, 204]. Lagrange [7, 146]. L’algèbre [218]. Lambert [24]. Landau [157]. language [208]. langue [208]. last [8, 86]. late [79, 37, 31]. Later [214]. Latin [212]. latine [212]. law [12]. Lazare [28]. Lebesgue [2]. Lectures [113, 218]. Leonard [40]. Leonardo [97]. Leopold [63, 121]. L’épistémologie [127]. letter [112]. letters [221, 26, 157]. lettre [112]. lettres [26]. Levi [109]. Levi-Civita [109]. Library [105]. liens [217]. life [43]. light [205]. lignes [93, 187]. linéaires [196]. linearum [155]. lines [93, 187]. link [156]. links [217]. literal [24]. littérales [24]. Little [171]. Lives [216]. livre [47, 36, 58]. logarithm [172]. Logarithmorum [215]. loi [12]. long [194]. Louis [214, 156]. Lucas [29, 68]. Lumières [134, 99, 181].

M [168, 108]. même [51]. machine [110]. Maclaurin [155]. Madrid [140]. maison [176]. manier [145]. manière [7]. manquant [156]. manuscript [105]. manuscripts [20]. manuscrit [105]. manuscrits [20]. marchande [106]. marchandes [192]. market [192]. Markov [9]. martingale [41]. Massau [82]. Math [37, 65]. mathematica [185]. Mathematical [214, 220, 205, 184, 211, 189, 25, 79, 208, 40, 106, 34, 116, 175, 139]. mathematician [43, 165]. Mathematicians [34, 198]. mathématicien [165]. mathématiciens [198]. Mathematics [64, 216, 32, 63, 1, 37, 53, 14, 71, 160, 59, 4, 13, 31, 60, 26, 196, 140, 25, 65, 157, 197, 176]. mathématique [63, 208, 106, 13, 116]. mathématiques [154, 32, 79, 53, 14, 4, 26, 196, 189, 176, 71, 190, 191, 220, 139]. Maurice [56]. Maurolico [47]. means [24]. measurement [215]. Mécanique [88, 8, 193]. mechanical [88]. mechanics [8, 67, 193, 88]. medieval [140, 58]. médiévale [58]. Meiji [208]. memoir [10, 57]. mémoire [10, 91, 57]. memory [91]. Mendes [192]. Menelaus [28]. Mésopotamie [168]. mesure [215]. metamathematical [110]. métamathématiques [110]. method [104, 57, 109, 126]. méthode [104, 57]. méthodes [23]. methods [11, 23]. Michael [187]. Middle [123]. Ministry [194]. Minkowski [143]. Minorsky [196]. missing [156]. mode [93]. model [30]. modèles [30]. modern [59, 209]. modernity [64]. mondiale [14]. Monograph [214]. morality [64]. Mordell [143]. Mosaics [71]. Mosaïques [71]. most [198]. motion [3]. Mouvance [174]. mouvement [3]. Moya [129, 203]. moyen [24, 123]. MR2582144 [154]. MS [140, 111]. Mullā [111]. multipas [23]. multiplication [54]. multistep [23]. multisubharmonic [5]. Mumbai [111]. Munitions [194]. Muslim [174]. musulmane [174].

Nacional [140]. Napier [172]. National [105]. nationale [105]. nature [79]. Naturphilosophie} [1]. naval [196]. navale [196]. Navier [91]. Nazi [34]. negative [48]. Netherlands [48]. network [193]. Netz [37]. neuere [113]. Newer [113]. Newton [155, 72]. Nicolas [196, 107]. Nicole [77]. nineteenth [67, 184, 6, 197]. nineteenth-century [67, 197]. Nippur [133]. no [37, 65]. Noether [17]. nombre [63]. nombres [168, 143, 87, 133, 187]. non [122, 18, 219, 196, 54, 123]. non-Archimedean [122]. non-continuous [123]. noninteger [18]. noninteger-order [18]. nonlinear [196]. notation [168, 169]. Notations [168]. note [37, 100, 129, 105]. notebooks [210]. notes [149]. notion [79]. nouvelle [208, 189]. Nouvelles [220]. nova [117]. novel [165]. Number [220, 63, 149, 143]. Numbers [222, 223, 224, 87, 68, 133, 187, 40]. numerical [169, 23, 168, 36]. numérique [36, 23].

ö [128]. observations [7]. occupation [154, 138, 217]. œuvre [87]. old [96, 117]. one [5]. only [110]. Opposition [157]. Opuscules [139]. order [156, 18, 92]. ordinaires [23]. ordinary [108, 23]. ordre [156, 92]. Oresme [77]. Organiser [217]. Organizing [217]. origin [23]. originalité [193]. originality [193]. origine [23]. origins [108]. oscillations [196]. other [69, 174, 37]. outil [106]. ouvrages [189]. ovales [62]. ovals [62]. overlapping [103]. overview [8].

painted [162]. Pappus [200]. parallel [172]. Parameterization [213]. paramétrage [213]. parler [51]. part [54, 151]. partage [174]. partial [139]. partie [151]. partielles [139]. partir [143]. Pascal [51]. Pasch [113]. Passer [144]. patterns [46]. Peano [113, 108, 211]. Pearson [194]. península [203, 140]. Pérez [129, 203]. Pergé [35]. periodic [193]. périodiques [193]. Perspective [30, 77]. Pétersbourg [41]. Petersburg [41]. Petite [171]. Peut [51]. Peut-on [51]. philology [197]. philosophie [4]. philosophy [4]. physical [205]. physics [186]. physique [186]. Pierre [44, 162]. planetary [7]. planètes [7]. plupart [198]. plurality [219]. plurisousharmoniques [5]. plus [151]. plusieurs [5]. Poincar [21]. Poincaré [173, 21, 52, 93, 193, 42]. Poinsot [156]. point [30, 3]. points [219]. polémique [52]. polynômes [57]. polynomials [57]. Poncelet [155]. portraits [163]. position [190, 191, 168]. Positive [48]. positivité [5]. positivity [5]. Post [138]. potential [11]. potential-theoretic [11]. Pour [32, 49, 208, 217, 196, 23]. Power [34]. powers [209]. Practical [216]. practice [143, 193]. practices [168, 184]. Pratica [192]. pratique [193]. pratiques [168, 143]. pre [172, 46]. pre-Classical [46]. pre-eminence [172]. précoce [105]. premier [199]. première [57]. premières [19, 106]. premiers [16, 217]. préparatoire [135]. Preparatory [135]. Prices [214]. principes [8]. Principios [129]. principles [8]. printed [106]. Priority [172]. probabilistic [173]. Probabilités [4, 20]. probabilities [20]. Probability [4]. probable [13]. problem [195, 200]. problème [195, 200]. Problems [109, 11]. programmatic [100]. projective [92]. proof [42]. proofs [11, 197]. proportion [123]. Proportionality [36]. proportionnalité [36]. proportions [123]. propos [190]. proposal [97]. Proposition [79]. Prospecting [46]. public [56]. Publier [138, 154].

Publish [154]. puis [176]. Puvis [162].

quadratic [212]. quadratiques [212]. quadrature [99, 188, 126]. Quantifier [133]. Quantify [133]. quantities [48]. quaternions [3, 3]. que [151]. quelques [21, 171, 198]. question [72]. questions [184, 77]. qu’on [112].

R.I.6 [111]. rôle [32, 200, 171]. racines [218]. Radical [209]. Randomness [214]. rapports [30, 215]. rate [78]. rational [8, 68]. Rationalité [58]. Rationality [58]. rationnelle [8]. reassessment [32, 185, 65]. recenseurs [217]. reception [48, 86, 113, 198, 123, 86, 113, 198, 123]. Recherches [7, 78]. récits [174]. reconstructions [25]. recouvrement [103]. recruitment [217]. recrutement [217]. rédaction [190]. réévaluation [32]. reflection [91]. reflections [168, 117, 37]. réflexion [91]. réflexions [168]. reform [79]. réforme [79]. règle [106]. règles [106]. relation [172]. relations [117]. relationship [30]. relationships [215]. relecture [58]. rentabilité [78]. rentes [78]. replay [58]. Researches [7, 78]. réseau [193]. resolution [139, 24, 139, 24]. résoudre [212]. Response [132]. restauration [212]. restituta} [117]. Restoration [212]. reticence [198, 198]. return [78]. Rev [37, 65]. Reviel [37]. review [60]. Revue [154, 190, 191]. revues [189]. Riemann [1, 185, 18]. rise [14]. role [32, 1, 200, 171, 65]. roles [97]. roman [165]. roots [218]. rule [106]. rules [106]. Russell [92, 52]. Ruy [192].

sa [77]. said [112]. Saint [41]. Saint-Pétersbourg [41]. sans [218]. Sanskrit [197]. Savoir [145]. school [122, 211, 9, 135]. Schooten [112]. Schubert [57]. Schur [150]. sciences [138]. scientific [91]. scientifiques [91]. se [110]. Second [14]. secondary [37]. seconde [14]. Self [163]. Self-portraits [163]. sell [176]. selon [187, 150]. sept [174]. series [179, 24, 77, 179, 24, 77]. Set [2]. seulement [110]. seules [7]. seventeen [174]. several [5]. shadow [163]. Shared [211]. sharing [174]. Shiite [174]. side [96]. siècle [134, 8, 86, 53, 105, 13, 6, 189, 198]. significant [198]. significatives [198]. similarity [123]. similitude [123]. Simon [187]. sixteenth [53]. Sketch [210]. so-called [42]. social [160]. sociétés [116]. societies [116]. solid [3]. solide [3]. solutions [193]. Solving [68, 212]. some [205, 21, 171, 96, 198, 116, 117]. source [140]. sources [203]. sous [138, 154, 217]. Space [204, 195]. spaces [16]. Spain [106]. Speculation [214]. spread [169]. Śrīdhara [180]. Stability [67]. stake [37]. States [14, 204]. statics [86]. statique [86]. Statistician [56]. statisticien [56]. statistics [20, 173]. statistique [20]. status [21, 52]. statut [21, 52]. Stevin [187, 198]. Stifel [187]. still [51]. stories [174]. String [116]. strings [179]. stroll [161]. structural [46]. struggles [34]. Students [220]. studies [25]. Study [220, 11, 40]. subterranea [216]. sujet [52]. surfaces [93]. Süss [34]. Sylvain [83]. Sylvester [68, 43]. Sylvestre [87]. Symbolic [146]. symbolique [146]. system [168, 169]. système [168].

T [57, 110]. tables [149, 7]. talk [51]. tapuscrits [195]. tardive [79]. taux [78]. Teachers [220]. Teaching [189, 208, 65]. Teixeira [221]. Telling [43]. tendancieux [30]. tendentious [30]. terms [110]. text [193]. textbook [48]. textes [79, 91, 21, 193]. Texts [214, 221, 79, 37, 91, 21, 31]. textual [25]. th [198]. their [122]. theorem [28, 103, 42]. théorème [28, 103]. theoretic [11]. théorie [10, 156, 29, 2, 30, 77, 6, 196, 215, 123, 218, 179]. Theory [220, 29, 222, 223, 224, 179, 11, 164, 10, 156, 149, 40, 2, 30, 77, 6, 196, 215, 123, 218]. thèse [128]. thesis [128]. thinking [48]. third [86, 68, 198]. Thomas [98]. Threat [132]. tiers [86]. time [160, 76]. tissus [71]. tool [106]. toujours [151]. tous [2]. tout [51, 151]. tradition [174]. traditional [116]. traditionnelles [116]. traité [35]. traités [171]. transcendents [164]. Tratado [203]. travaux [198]. treatise [35]. treatises [171]. triangle [51, 189]. trigonometric [179]. trigonométriques [179]. troisième [198]. Turing [128]. turn [219]. two [20, 100, 76]. types [100]. typescripts [195].

uncle [162]. Unis [14, 204]. United [14, 204]. Universal} [92]. unprecedented [10]. unpublished [63, 20]. unwritten [54]. Ur [76]. usages [204, 133]. Use [204, 105]. useful [71]. uses [133]. utiles [71]. utilisation [105].

valeur [127]. value [11, 127]. Vandermonde [175]. variable [5]. variables [5]. Variations [28]. vector [16]. vectoriels [16]. vendre [176]. version [111]. versus} [197]. vibrantes [179]. vibrating [179]. Viète [117]. view [30]. VII [36]. visualizations [219]. vocables [110]. Vorlesungen [113, 218]. vue [30].

Wallis [72, 88]. war [34, 14]. Weyl [195, 186, 199]. where [117]. which [112]. Whitehead [92]. Wilhelm [34]. Wilson [157]. without [218]. Witt [42]. Wolfgang [118]. Work [214, 87, 198]. World [14]. writing [91, 93]. writings [145, 88].

X [58]. XIXe [86, 6, 189]. XIXth [86]. XVe [105]. XVI [198]. XVI-th [198]. XVIe [53]. XVIIe [13].

years [25].

Zentralblatt [217].

References

Bottazzini:1995:NRR

[1] Umberto Bottazzini and Rossana Tazzioli. Naturphilosophie and its role in Riemann’s mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(1):3–38, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/1/html/smf_rhm_1_3-38.html.

Gispert:1995:TEF

[2] Hélène Gispert. La théorie des ensembles en France avant la crise de 1905: Baire, Borel, Lebesgue … et tous les autres. (French) [Set theory in France before the crisis of 1905: Baire, Borel, Lebesgue … and all the others]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(1):39–81, 1995. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/1/html/smf_rhm_1_39-81.html.

Sinegre:1995:QMD

[3] Luc Sinègre. Les quaternions et le mouvement du solide autour d’un point fixe chez Hamilton. (French) [Quaternions and the motion of the solid around a fixed point at Hamilton]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(1):83–109, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/1/html/smf_rhm_
1_83-109.html.

Martin:1995:PPM

[4] Thierry Martin. Probabilités et philosophie des mathématiques chez Cournot. (French) [Probability and philosophy of mathematics at Cournot]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(1):111–138, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/1/html/smf_rhm_1_111-138.html.

Lelong:1995:DVP

[5] Pierre Lelong. D’une variable à plusieurs variables en analyse complexe: les fonctions plurisousharmoniques et la positivité (1942–1962). (French) [From one variable to several variables in complex analysis: multisubharmonic functions and positivity (1942–1962)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(1):139–157, 1995. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/1/html/smf_rhm_1_139-157.html.

Nabonnand:1995:CHT

[6] Philippe Nabonnand. Contribution à l’histoire de la théorie des géodésiques au XIXe siècle. (French) [Contribution to the history of geodesic theory in the nineteenth century]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(2):159–200, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/1/
html/smf_rhm_1_159-200.html.

Galuzzi:1995:LER

[7] Massimo Galuzzi. Lagrange’s essay “Recherches sur la manière de former des tables des planètes d’après les seules observations”. (French) [Lagrange’s essay “researches” on how to form planetary tables based on observations alone”]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(2):201–233, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/1/html/smf_rhm_1_201-233.html.

Chatzis:1995:ADP

[8] Konstantinos Chatzis. Un aperçu de la discussion sur les principes de la mécanique rationnelle en France à la fin du siècle dernier. (French) [An overview of the discussion on the principles of rational mechanics in France at the end of the last century]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(2):235–270, 1995. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/1/html/smf_rhm_1_235-270.html.

Margenstern:1995:LCM

[9] Maurice Margenstern. L’école constructive de Markov. (French) [The constructive school of Markov]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 1(2):271–305, 1995. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/1/html/smf_rhm_
1_271-305.html.

Belhoste:1996:AMI

[10] Bruno Belhoste. Autour d’un mémoire inédit: la contribution d’Hermite au développement de la théorie des fonctions elliptiques. (French) [Around an unprecedented memoir: Hermite’s contribution to the development of the theory of elliptic functions]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(1):1–66, 1996. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/2/html/smf_rhm_
2_1-66.html.

Archibald:1996:ATE

[11] Thomas Archibald. From attraction theory to existence proofs: the evolution of potential-theoretic methods in the study of boundary-value problems, 1860–1890. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(1):67–93, 1996. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/2/html/smf_rhm_2_67-93.html.

Vilain:1996:LGD

[12] Christiane Vilain. La loi galiléenne et la dynamique de Huygens. (French) [Galilean law and the dynamics of Huygens]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(1):95–117, 1996. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/2/
html/smf_rhm_2_95-117.html.

Meusnier:1996:LMD

[13] Norbert Meusnier. L’émergence d’une mathématique du probable au XVIIe siècle. (French) [The emergence of a mathematics of the probable in the 17th century]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(1):119–147, 1996. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/2/html/smf_rhm_2_119-147.html.

Dalmedico:1996:LMA

[14] Amy Dahan Dalmedico. L’essor des mathématiques appliquées aux États-Unis: l’impact de la seconde guerre mondiale. (French) [The rise of applied mathematics in the United States: the impact of the Second World War]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(2):149–213, 1996. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/2/html/smf_rhm_2_149-213.html.

DellAglio:1996:GCC

[15] Luca Dell’Aglio. On the genesis of the concept of covariant differentiation. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(2):215–264, 1996. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/2/html/smf_rhm_2_215-264.html.

Dorier:1996:GPE

[16] Jean-Luc Dorier. Genèse des premiers espaces vectoriels de fonctions. (French) [Genesis of the first vector spaces of functions]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 2(2):265–307, 1996. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/2/html/smf_rhm_2_265-307.html.

Gray:1997:AGB

[17] Jeremy Gray. Algebraic geometry between Noether and Noether — a forgotten chapter in the history of algebraic geometry. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 3(1):1–48, 1997. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/3/html/smf_rhm_3_1-48.html.

Dugowson:1997:LPR

[18] Stéphane Dugowson. L’élaboration par Riemann d’une définition de la dérivation d’ordre non entier. (French) [The Riemann elaboration of a definition of the noninteger-order derivation]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 3(1):49–97, 1997. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/3/html/smf_rhm_3_49-97.html.

Annaratone:1997:PDF

[19] Silvia Annaratone. Les premières démonstrations de la formule intégrale de Fourier. (French) [The first demonstrations of the integral formula of Fourier]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 3(1):99–136, 1997. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/3/html/smf_rhm_3_99-136.html.

Bru:1997:SCP

[20] Bernard Bru, Marie-France Bru, and Olivier Bienaymé. La statistique critiquée par le calcul des probabilités: deux manuscrits inédits d’Irenée Jules Bienaymé. (French) [The statistics criticized by the calculation of the probabilities: two unpublished manuscripts of Irenée Jules Bienaymé]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 3(2):137–239, 1997. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/3/html/smf_rhm_3_137-239.html.

Herreman:1997:SGQ

[21] Alain Herreman. Le statut de la géométrie dans quelques textes sur l’homologie, de Poincaré aux années 1930. (French) [The status of geometry in some texts on homology, from Poincar to the 1930s]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 3(2):241–293, 1997. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/3/html/smf_rhm_3_241-293.html.

Redaction-en-chef:1998:Ea

[22] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(1):1–4, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/4/html/smf_rhm_
4_1-4.html.

Tournes:1998:LMM

[23] Dominique Tournès. L’origine des méthodes multipas pour l’intégration numérique des équations différentielles ordinaires. (French) [The origin of multistep methods for the numerical integration of ordinary differential equations]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(1):5–72, 1998. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/4/html/smf_rhm_4_5-72.html.

Lubet:1998:LCR

[24] Jean-Pierre Lubet. De Lambert à Cauchy: la résolution des équations littérales par le moyen des séries. (French) [From Lambert to Cauchy: the resolution of literal equations by means of series]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(1):73–129, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/4/
html/smf_rhm_4_73-129.html.

Saito:1998:MRT

[25] Ken Saito. Mathematical reconstructions out, textual studies in: 30 years in the historiography of Greek mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(1):131–142, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/4/
html/smf_rhm_4_131-142.html.

Peiffer:1998:FMP

[26] Jeanne Peiffer. Faire des mathématiques par lettres. (French) [Doing mathematics by letters]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(1):143–157, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/4/html/smf_rhm_4_143-157.html.

Redaction-en-chef:1998:Eb

[27] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):159–161, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/4/html/smf_rhm_
4_159-161.html.

Chemla:1998:LCG

[28] Karine Chemla. Lazare Carnot et la généralité en géométrie. Variations sur le théorème dit de Menelaus. (French) [Lazare Carnot and generality in geometry. Variations on the theorem of Menelaus]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):163–190, 1998. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/4/html/smf_rhm_4_163-190.html.

Decaillot:1998:LEL

[29] Anne-Marie Décaillot. L’arithméticien Édouard Lucas (1842–1891): théorie et instrumentation. (French) [The arithmetician Édouard Lucas (1842–1891): Theory and instrumentation]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):191–236, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/4/
html/smf_rhm_4_191-236.html.

Lascar:1998:PHR

[30] Daniel Lascar. Perspective historique sur les rapports entre la théorie des modèles et l’algèbre. Un point de vue tendancieux. (French) [Historical perspective on the relationship between model theory and algebra. A tendentious point of view]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):237–260, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/4/html/smf_rhm_4_237-260.html.

Netz:1998:DTL

[31] Reviel Netz. Deuteronomic texts: late antiquity and the history of mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):261–288, 1998. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/

RevueHistoireMath/4/html/smf_rhm_4_261-288.html. See commen-

taries [37].

Belhoste:1998:PRD

[32] Bruno Belhoste. Pour une réévaluation du rôle de l’enseignement dans l’histoire des mathématiques. (French) [For a reassessment of the role of education in the history of mathematics]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 4(2):289–304, 1998. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/4/
html/smf_rhm_4_289-304.html. See comment [65].

Redaction-en-chef:1999:Ea

[33] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):1–5, 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/5/html/smf_rhm_
5_1-5.html.

Remmert:1999:MWP

[34] Volker R. Remmert. Mathematicians at war. Power struggles in Nazi Germany’s mathematical community: Gustav Doetsch and Wilhelm Süss. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):7–59, 1999. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/5/html/smf_rhm_5_7-59.html.

Decorps-Foulquier:1999:FDT

[35] Micheline Decorps-Foulquier. Sur les figures du traité des Coniques d’Apollonios de Pergé édité par Eutocius d’Ascalon. (French) [On the figures of the treatise of Conics of Apollonius of Pergé edited by Eutocius of Ascalon]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):61–82, 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/5/html/smf_rhm_5_61-82.html.

Rommevaux:1999:PNL

[36] Sabine Rommevaux. La proportionnalité numérique dans le Livre VII des Éléments de Campanus. (French) [Numerical proportionality in Book VII of Elements of Campanus]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):83–126, 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/5/html/smf_rhm_5_83-126.html.

Chemla:1999:CEO

[37] Karine Chemla. Commentaries, editions and other secondary texts: what is the stake for the history of mathematics? Reflections inspired by Reviel Netz’s note: “Deuteronomic texts: late antiquity and the history of mathematics” [Rev. Histoire Math. 4 (1998), no. 3, 261–288; (2000g:01012)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):127–148, 1999. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/5/html/smf_rhm_5_127-148.html. See [31].

Fowler:1999:II

[38] David Fowler. Inventive interpretations. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(1):149–153, 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/5/
html/smf_rhm_5_149-153.html.

Redaction-en-chef:1999:Eb

[39] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2):155–157, 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/5/html/smf_rhm_
5_155-157.html.

Fenster:1999:LDH

[40] Della D. Fenster. Leonard Dickson’s History of the theory of numbers: an historical study with mathematical implications. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2): 159–179 (2000), 1999. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/5/html/smf_rhm_5_159-179.html.

Bru:1999:BMS

[41] Bernard Bru, Marie-France Bru, and Kai Lai Chung. Borel et la martingale de Saint-Pétersbourg. (French) [Borel and the martingale of Saint Petersburg]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2):181–247 (2000), 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/5/html/smf_rhm_5_181-247.html.

Ton-That:1999:PPC

[42] Tuong Ton-That and Thai-Duong Tran. Poincaré’s proof of the so-called Birkhoff–Witt theorem. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2):249–284 (2000), 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/5/html/smf_rhm_
5_249-284.html.

Parshall:1999:TLM

[43] Karen Hunger Parshall. Telling the life of a mathematician: the case of J. J. Sylvester. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2):285–302 (2000), 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/5/html/smf_rhm_5_285-302.html.

Anonymous:1999:PD

[44] Anonymous. Pierre Dugac. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 5(2):303–315 (2000), 1999. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/5/html/smf_rhm_
5_303-315.html.

Redaction-en-chef:2000:Ea

[45] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(1):1–3, 2000. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/6/html/smf_rhm_
6_1-3.html.

Hoyrup:2000:GPP

[46] Jens Høyrup. Geometrical patterns in the pre-Classical Greek area. Prospecting the borderland between decoration, art, and structural inquiry. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(1):5–58, 2000. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/6/html/smf_rhm_6_5-58.html.

Sutto:2000:CDL

[47] Jean-Pierre Sutto. Le compendium du 5e livre des Éléments d’Euclide de Francesco Maurolico. (French) [The compendium of the 5th book of euclid’s Elements by Francesco Maurolico]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(1):59–94, 2000. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/6/
html/smf_rhm_6_59-94.html.

Beckers:2000:PTC

[48] Danny J. Beckers. Positive thinking. Conceptions of negative quantities in the Netherlands and the reception of Lacroix’s algebra textbook. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(1):95–126, 2000. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/6/html/smf_rhm_6_95-126.html.

Tournes:2000:PHD

[49] Dominique Tournès. Pour une histoire du calcul graphique. (French) [For a history of graphing]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(1):127–161, 2000. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/6/html/smf_rhm_6_127-161.html.

Redaction-en-chef:2000:Eb

[50] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(2):163–166, 2000. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/6/html/smf_rhm_
6_1-3.html.

Kyriacopoulos:2000:PTM

[51] Laurent Kyriacopoulos. Peut-on tout de même parler d’un “triangle de Pascal”?. (French) [Can we still talk about a “triangle of Pascal”?]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(2):167–217, 2000. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/6/html/smf_rhm_6_167-217.html.

Nabonnand:2000:PEP

[52] Philippe Nabonnand. La polémique entre Poincaré et Russell au sujet du statut des axiomes de la géométrie. (French) [The controversy between Poincaré and Russell about the status of axioms of geometry]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(2):219–269, 2000. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/6/html/smf_rhm_6_219-269.html.

Compere:2000:CIE

[53] Marie-Madeleine Compère. Les cadres institutionnels de l’enseignement des mathématiques au XVIe siècle. (French) [Institutional frameworks of mathematics education in the sixteenth century]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(2):271–292, 2000. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/6/html/smf_rhm_6_271-292.html.

Proust:2000:MBP

[54] Christine Proust. La multiplication babylonienne: la part non écrite du calcul. (French) [Babylonian multiplication: the unwritten part of the calculation]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 6(2):293–303, 2000. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/6/html/smf_rhm_6_293-303.html.

Redaction-en-chef:2001:Ea

[55] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):1–4, 2001. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/7/pdf/smf_rhm_7_1-4.
pdf.

Armatte:2001:MFS

[56] Michel Armatte. Maurice Fréchet statisticien, enquêteur et agitateur public. (French) [Maurice Fréchet: Statistician, investigator and public agitator]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):7–65, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_7-65.html.

Mignotte:2001:PMG

[57] Maurice Mignotte and Doru Ştefănescu. La première méthode générale de factorisation des polynômes. Autour d’un mémoire de F. T. Schubert. (French) [The first general method for factoring polynomials. On a memoir of F. T. Schubert]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):67–89, 2001. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/7/html/smf_rhm_7_67-89.html.

Rommevaux:2001:RER

[58] Sabine Rommevaux. Rationalité, exprimabilité: une relecture médiévale du Livre X des Éléments d’Euclide. (French) [Rationality, expressability: a medieval replay of Book X of Elements of Euclid]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):91–119, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_91-119.html.

Knorr:2001:IMM

[59] Wilbur R. Knorr. The impact of modern mathematics on ancient mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):121–135, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_121-135.html.

Ogawa:2001:RHJ

[60] Tsukane Ogawa. A review of the history of Japanese mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(1):137–155, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_137-155.html.

Redaction-en-chef:2001:Eb

[61] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(2):157–159, 2001. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http:/

/smf4.emath.fr/Publications/RevueHistoireMath/7/pdf/smf_rhm_
7_157-159.pdf.

Barbin:2001:AFE

[62] Évelyne Barbin and René Guitart. Algèbre des fonctions elliptiques et géométrie des ovales cartésiennes. (French) [Algebra of elliptic functions and geometry of cartesian ovals]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(2):161–205, 2001. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/7/html/smf_rhm_
7_161-205.html.

Boniface:2001:CNM

[63] Jacqueline Boniface and Norbert Schappacher. “Sur le concept de nombre en mathématique”: cours inédit de Leopold Kronecker à Berlin (1891). (French) [“ On the concept of number in mathematics’: unpublished course of Leopold Kronecker in Berlin (1891)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(2):206–275, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_207-275.html.

Dear:2001:MMC

[64] Peter Dear. Mathematics and morality on the cusp of modernity. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(2):277–293, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_277-293.html.

Schubring:2001:CTR

[65] Gert Schubring. Comment: “Toward a reassessment of the role of teaching in the history of mathematics” [Rev. Histoire Math. 4 (1998), no. 2, 289–304; (2001a:01075)] by B. Belhoste. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 7(2):295–305, 2001. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/7/html/smf_rhm_7_295-305.html. See [32].

Redaction-en-chef:2002:Ea

[66] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(1):1–4, 2002. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/8/pdf/smf_rhm_8_1-4.
pdf.

Darrigol:2002:SIN

[67] Olivier Darrigol. Stability and instability in nineteenth-century fluid mechanics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(1):5–65, 2002. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/8/html/smf_rhm_8_5-65.html.

Lavrinenko:2002:SIT

[68] Tatiana Lavrinenko. Solving an indeterminate third degree equation in rational numbers. Sylvester and Lucas. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(1):67–111, 2002. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/8/html/smf_rhm_
8_67-111.html.

Ageron:2002:LAC

[69] Pierre Ageron. L’autre axiome de choix. (French) [The other axiom of choice]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(1):113–140, 2002. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/8/html/smf_rhm_8_113-140.html.

Redaction-en-chef:2002:Eb

[70] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(2):141–144, 2002. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http:/

/smf4.emath.fr/Publications/RevueHistoireMath/8/pdf/smf_rhm_
8_141-144.pdf.

Decaillot:2002:GTM

[71] Anne-Marie Décaillot. Géométrie des tissus. Mosaïques. Échiquiers. Mathématiques curieuses et utiles. (French) [Geometry of cloth. Mosaics. Chess. Curious and useful mathematics]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(2):145–206, 2002. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/8/
html/smf_rhm_8_145-206.html.

Loget:2002:WEH

[72] François Loget. Wallis entre Hobbes et Newton. La question de l’angle de contact chez les Anglais. (French) [Wallis between Hobbes and Newton. The question of contact angle in English]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(2):207–262, 2002. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/8/html/smf_rhm_8_207-262.html.

Netz:2002:TC

[73] Reviel Netz. It’s not that they couldn’t. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 8(2):263–289, 2002. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/8/
html/smf_rhm_8_263-289.html.

Redaction-en-chef:2003:DDF

[74] La Rédaction en chef. Éditorial du fascicule 1. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):1–4, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/9/
pdf/smf_rhm_9_1-8.pdf.

Editors-in-Chief:2003:Ea

[75] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):5–8, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/9/pdf/smf_rhm_9_1-8.
pdf.

Quillien:2003:DCE

[76] Jacques Quillien. Deux cadastres de l’époque d’Ur III. (French) [Two cadastres from the time of Ur III]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):9–31, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/9/html/smf_rhm_
9_9-31.html.

Mazet:2003:TSN

[77] Edmond Mazet. La théorie des séries de Nicole Oresme dans sa perspective aristotélicienne. Questions 1 et 2 sur la géométrie d’Euclide. (French) [The series theory of Nicole Oresme in its Aristotelian perspective. Questions 1 and 2 about the geometry of Euclid]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):33–80, 2003. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/9/html/smf_rhm_9_33-80.html.

Biondi:2003:RRD

[78] Yuri Biondi. Les Recherches sur les rentes de Duvillard (1787) et le taux interne de rentabilité. (French) [The Annuities Researches of Duvillard (1787) and the internal rate of return]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):81–130, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/9/
html/smf_rhm_9_81-130.html.

Bernard:2003:CDN

[79] Alain Bernard. Comment définir la nature des textes mathématiques de l’antiquité grecque tardive? Proposition de réforme de la notion de “textes deutéronomiques”. (French) [How to define the nature of the mathematical texts of late Greek antiquity? Proposition of reform of the concept of “deuteronomic texts”]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(1):131–173, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/9/html/smf_rhm_9_131-173.html.

Redaction-en-chef:2003:E

[80] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(2):175–177, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http:/

/smf4.emath.fr/Publications/RevueHistoireMath/9/pdf/smf_rhm_
9_175-180.pdf.

Editors-in-Chief:2003:Eb

[81] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(2):178–180, 2003. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http:/

/smf4.emath.fr/Publications/RevueHistoireMath/9/pdf/smf_rhm_
9_175-180.pdf.

Tournes:2003:JMI

[82] Dominique Tournès. Junius Massau et l’intégration graphique. (French) [Junius Massau and graphical integration]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(2): 181–252 (2004), 2003. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/9/html/smf_rhm_9_181-252.html.

Raina:2003:BJE

[83] Dhruv Raina. Betwixt Jesuit and enlightenment historiography: Jean-Sylvain Bailly’s history of Indian astronomy. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 9(2): 253–306 (2004), 2003. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/9/html/smf_rhm_9_253-306.html.

Redaction-en-chef:2004:Ea

[84] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(1):1–3, 2004. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/10/pdf/smf_rhm_10_1-
6.pdf.

Editors-in-Chief:2004:Ea

[85] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(1):4–6, 2004. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/10/pdf/smf_rhm_10_1-
6.pdf.

Chatzis:2004:RSG

[86] Konstantinos Chatzis. La réception de la statique graphique en France durant le dernier tiers du XIXe siècle. (French) [The reception of graphic statics in France during the last third of the XIXth century]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(1):7–43, 2004. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/10/html/smf_rhm_10_7-43.html.

Lamande:2004:CNF

[87] Pierre Lamandé. La conception des nombres en France autour de 1800: l’œuvre didactique de Sylvestre François Lacroix. (French) [The design of numbers in France around 1800: the didactic work of Sylvestre François Lacroix]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(1):45–106, 2004. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/10/html/smf_rhm_10_45-106.html.

Maieru:2004:MGE

[88] Luigi Maierù. Mécanique et géométrie dans les écrits de mécanique de John Wallis. Le calcul du centre de gravité. (French) [Mechanics and geometry in the mechanical writings of John Wallis. The calculation of the center of gravity]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(1):107–134, 2004. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/10/html/
smf_rhm_10_107-134.html.

Redaction-en-chef:2004:Eb

[89] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(2):135–137, 2004. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/10/html/
smf_rhm_10_135-140.html. À la mémoire de René Taton.

Editors-in-Chief:2004:Eb

[90] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(2):138–140, 2004. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/10/pdf/smf_rhm_
10_135-140.pdf.

Graber:2004:AOM

[91] Frédéric Graber. Article ou mémoire? Une réflexion comparative sur l’écriture des textes scientifiques. Navier et l’écoulement des fluides (1822– 1827). (French) [Article or memory? A comparative reflection on the writing of scientific texts. Navier and fluid flow (1822–1827)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(2):141–185 (2005), 2004. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/10/html/smf_rhm_10_141-185.html.

Gandon:2004:RUA

[92] Sébastien Gandon. Russell et l’Universal algebra de Whitehead: la géométrie projective entre ordre et incidence (1898–1903). (French) [Russell and Whitehead’s Universal algebra: the projective geometry between order and incidence (1898–1903)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(2): 187–256 (2005), 2004. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/10/html/smf_rhm_10_187-256.html.

Robadey:2004:EME

[93] Anne Robadey. Exploration d’un mode d’écriture de la généralité: l’article de Poincaré sur les lignes géodésiques des surfaces convexes (1905). (French) [Exploration of a mode of writing of generality: the article of Poincaré on the geodesic lines of convex surfaces (1905)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 10(2):257–318 (2005), 2004. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/10/html/smf_rhm_10_257-318.html.

Redaction-en-chef:2005:Ea

[94] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):1–3, 2005. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_1-6.html;   http://smf4.emath.fr/Publications/RevueHistoireMath/
11/pdf/smf_rhm_11_1-6.pdf.

Editors-in-Chief:2005:Ea

[95] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):4–6, 2005. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_1-6.html;   http://smf4.emath.fr/Publications/RevueHistoireMath/
11/pdf/smf_rhm_11_1-6.pdf.

Melville:2005:ASA

[96] Duncan J. Melville. The area and the side I added: some old Babylonian geometry. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):7–21, 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_7-21.html.

Hoyrup:2005:LFA

[97] Jens Høyrup. Leonardo Fibonacci and Abbaco culture. A proposal to invert the roles. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):23–56, 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_23-56.html.

Maclean:2005:THC

[98] Ian Maclean. Thomas Harriot on combinations. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):57–88, 2005. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/11/
html/smf_rhm_11_57-88.html.

Jacob:2005:IQD

[99] Marie Jacob. Interdire la quadrature du cercle à l’Académie: une décision autoritaire des Lumières?. (French) [To forbid the quadrature of the circle at the Academy: an authoritarian decision of the Enlightened?]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):89–139, 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_89-139.html.

Netz:2005:PNT

[100] Reviel Netz. A programmatic note: on two types of intertextuality. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):143–155, 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_143-155.html.

Redaction-en-chef:2005:Eb

[101] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):157–159, 2005. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_157-162.html.

Editors-in-Chief:2005:Eb

[102] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):160–162, 2005. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_157-162.html.

Maurey:2005:GDT

[103] Bernard Maurey and Jean-Pierre Tacchi. La genèse du théorème de recouvrement de Borel. (French) [The genesis of Borel’s overlapping theorem]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):163–204 (2006), 2005. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_163-204.html.

Brezinski:2005:MC

[104] Claude Brezinski. La méthode de Cholesky. (French) [The method of Cholesky]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):205–238 (2006), 2005. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_205-238.html.

Lamasse:2005:UPA

[105] Stéphane Lamassé. Une utilisation précoce de l’algèbre en France au XVe siècle. Note sur le manuscrit 1339 de la Bibliothèque nationale de France. (French) [Early use of algebra in France in the fifteenth century. Note on manuscript 1339 of the National Library of France]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):239–255 (2006), 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_239-255.html.

Labarthe:2005:RCP

[106] Marie-Hélène Labarthe. Les règles de compagnie, dans les premières arithmétiques imprimées des Espagnes: de la règle marchande à l’outil mathématique. (French) [The rules of company, in the first printed arithmetic of Spain: from the commercial rule to the mathematical tool]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(2):257–313 (2006), 2005. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/11/html/smf_rhm_11_257-313.html.

Spiesser:2006:LNC

[107] Maryvonne Spiesser. L’algèbre de Nicolas Chuquet dans le contexte français de l’arithmétique commerciale. (French) [The algebra of Nicolas Chuquet in the French context of commercial arithmetic]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(1):7–33 (2007), 2006. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/12/html/smf_rhm_12_7-33.html.

Luciano:2006:OFA

[108] Erika Luciano. At the origins of functional analysis: G. Peano and M. Gramegna on ordinary differential equations. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12 (1):35–79 (2007), 2006. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/12/html/smf_rhm_12_35-79.html.

Nastasi:2006:PML

[109] Pietro Nastasi and Rossana Tazzioli. Problems of method in Levi-Civita’s contributions to hydrodynamics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(1):81–118 (2007), 2006. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/12/html/
smf_rhm_12_81-118.html.

Kromer:2006:MGS

[110] Ralf Krömer. La “machine de Grothendieck” se fonde-t-elle seulement sur des vocables métamathématiques? Bourbaki et les catégories au cours des années cinquante. (French) [Is the ’Grothendieck’ machine based on metamathematical terms only? Bourbaki and categories in the fifties]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(1):119–162 (2007), 2006. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/12/html/smf_rhm_12_119-162.html.

Brentjes:2006:ENA

[111] Sonja Brentjes. An exciting new Arabic version of Euclid’s Elements: Ms Mumbai, Mullā Fīrūz R.I.6. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(2):169–197 (2007), 2006. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/12/html/
smf_rhm_12_169-197.html.

Maronne:2006:LDS

[112] Sébastien Maronne. Sur une lettre de Descartes à Schooten qu’on dit de 1639. (French) [On a letter from Descartes to Schooten which is said to be 1639]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(2):199–248 (2007), 2006. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/12/html/

smf_rhm_12_199-248.html. Correspondence: Descartes, René–van Schooten, Frans.

Gandon:2006:RVN

[113] Sébastien Gandon. La réception des Vorlesungen über neuere Geometrie de Pasch par Peano. (French) [The reception of Pasch’s Lectures on Newer Geometry by Peano]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 12(2):249–290 (2007), 2006. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/12/html/
smf_rhm_12_249-290.html.

Redaction-en-chef:2007:Ea

[114] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(1):1–3, 2007. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/13/rhm13-1-edito.pdf.

Editors-in-Chief:2007:Ea

[115] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(1):4–6, 2007. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/Publications/RevueHistoireMath/13/rhm13-1-edito.pdf.

Vandendriessche:2007:JFA

[116] Éric Vandendriessche. Les jeux de ficelle: une activité mathématique dans certaines sociétés traditionnelles. (French) [String games: a mathematical activity in some traditional societies]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(1):7–84, 2007. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/13/
html/smf_rhm_13_7-84.html.

Panza:2007:WNW

[117] Marco Panza. What is new and what is old in Viète’s analysis restituta and algebra nova, and where do they come from? Some reflections on the relations between algebra and analysis before Viète. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(1):85–153, 2007. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/13/html/smf_rhm_13_85-153.html.

Mazliak:2007:EBW

[118] Laurent Mazliak. On the exchanges between Wolfgang Doeblin and Bohuslav Hostinský. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(1):155–180, 2007. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/13/html/
smf_rhm_13_155-180.html.

Redaction-en-chef:2007:Eb

[119] La Rédaction en chef. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(2):181–183, 2007. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/13/pdf/rhm13-2-
edito.pdf.

Editors-in-Chief:2007:Eb

[120] The Editors-in-Chief. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(2):184–186, 2007. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/13/pdf/rhm13-2-
edito.pdf.

Brechenmacher:2007:CEC

[121] Frédéric Brechenmacher. La controverse de 1874 entre Camille Jordan et Leopold Kronecker. (French) [The controversy of 1874 between Camille Jordan and Leopold Kronecker]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(2):187–257, 2007. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications/RevueHistoireMath/13/html/
smf_rhm_13_187-257.html.

Cerroni:2007:CHD

[122] Cinzia Cerroni. The contributions of Hilbert and Dehn to non-Archimedean geometries and their impact on the Italian school. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(2):259–299, 2007. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/13/html/smf_rhm_13_259-299.html.

Rommevaux:2007:SED

[123] Sabine Rommevaux. La similitude des équimultiples dans la définition de la proportion non continue de l’édition des Éléments d’Euclide par Campanus: une difficulté dans la réception de la théorie des proportions au Moyen Âge. (French) [The similarity of equimultiples in the definition of the non-continuous proportion of the edition of Euclid’s Elements by Campanus: a difficulty in the reception of the theory of proportions in the Middle Ages]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 13(2):301–322, 2007. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/13/html/smf_rhm_13_301-322.html.

Peiffer:2008:Ea

[124] Jeanne Peiffer. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):1–2, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/14/pdf/rhm14-1-edito.pdf.

Peiffer:2008:Eb

[125] Jeanne Peiffer. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):3–4, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/14/pdf/rhm14-1-edito.pdf.

Paradis:2008:FMQ

[126] Jaume Paradís, Josep Pla, and Pelegrí Viader. Fermat’s method of quadrature. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):5–51, 2008. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/14/pdf/smf_rhm_14_5-51.pdf.

Barberousse:2008:VCA

[127] Anouk Barberousse. La valeur de la connaissance approchée. L’épistémologie de l’approximation d’Émile Borel. (French) [The value of approximate knowledge. The epistemology of Émile Borel’s approximation]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):53–75, 2008. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/14/pdf/smf_rhm_14_53-75.pdf.

Cassou-Nogues:2008:GT

[128] Pierre Cassou-Noguès. Gödel et la thèse de Turing. (French) [G ö del and the thesis of Turing]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):77–111, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/14/pdf/smf_rhm_14_77-111.pdf.

Silva:2008:NPM

[129] M. Céu Silva and Antoni Malet. A note on Pérez de Moya’s Principios de geometria (1584). Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(1):113–133, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/14/pdf/smf_rhm_14_113-133.pdf.

Schappacher:2008:Ea

[130] Norbert Schappacher. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):135–136, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/RevueHistoireMath/14/pdf/rhm14-2-edito.pdf.

Schappacher:2008:Eb

[131] Norbert Schappacher. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):137–138, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/RevueHistoireMath/14/pdf/rhm14-2-edito.pdf.

Andersen:2008:JUT

[132] Hanne Andersen (Centaurus), Roger Ariew, Moti Feingold (Perspectives on Science), A. K. Bag (Indian Journal of History of Science), June Barrow-Green, Benno van Dalen (Historia mathematica), Keith Benson (History, Philosophy of the Life Sciences), Marco Beretta (Nuncius), Michel Blay (Revue d’Histoire des Sciences), Cornelius Borck (Berichte zür Wissenschaftsgeschichte), Geof Bowker, Technology Susan Leigh Star (Science, Human Values), Massimo Bucciantini, Michele Camerota (Galilaeana: Journal of Galilean Studies), Jed Buchwald, Jeremy Gray (Archive for History of Exact Sciences), Vincenzo Cappelletti, Guido Cimino (Physis), Mark Clark, Alex Keller (ICON), Roger Cline (International Journal for the History of Engineering & Technology), Stephen Clucas, Stephen Gaukroger (Intellectual History Review), Hal Cook, Anne Hardy (Medical History), Alexandre Métraux Leo Corry, Jürgen Renn (Science in Context), Brian Dolan, Bill Luckin (Social History of Medicine), Hilmar Duerbeck, Wayne Orchiston (Journal of Astronomical History & Heritage), Hans-Jörg Rheinberger Moritz Epple, Mikael Hård, Technik und Medizin) Volker Roelcke (NTM: Zeitschrift für Geschichte der Wissenschaften, Steven French (Metascience), Paul Farber (Journal of the History of Biology), Mary Fissell, Randall Packard (Bulletin of the History of Medicine), Robert Fox (Notes, Records of the Royal Society), Jim Good (History of the Human Sciences), Willem Hackmann (Bulletin of the Scientific Instrument Society), Robert Halleux (Archives Internationales d’Histoire des Sciences), Bosse Holmqvist (Lychnos), Michael Hoskin (Journal for the History of Astronomy), Ian Inkster (History of Technology), Marina Frasca Spada (Studies in History, Philosophy of Science), Nick Jardine (Studies in History, Philosophy of Biological, Biomedical Sciences), Trevor Levere (Annals of Science), Bernard Lightman (Isis), Christoph Lüthy (Early Science, Medicine), Michael Lynch (Social Studies of Science), Stephen McCluskey, Clive Ruggles (Archaeostronomy: the Journal of Astronomy in Culture), Peter Morris (Ambix), E. Charles Nelson (Archives of Natural History), Ian Nicholson (Journal of the History of the Behavioural Sciences), Kathy Olesko (Osiris), Liliane Pérez (Documents pour l’Histoire des Techniques), Iwan Rhys Morus (History of Science), John Rigden, Roger H. Stuewer (Physics in Perspective), Julio Samsó (Suhayl: Journal for the History of the Exact, Natural Sciences in Islamic Civilisation), Simon Schaffer (British Journal for the History of Science), Norbert Schappacher (Revue d’histoire des mathématiques), Claire Strom (Agricultural History), Paul Unschuld (Sudhoffs Archiv), Peter Weingart (Minerva), Stefan Zamecki (Kwartalnik Historii Nauki i Techniki), and Huib Zuidervaart (Studium. Tijdschrift Voor Wetenschaps en Universiteitsgeschiedenis/Revue de l’Histoire des Sciences et des Universités). Journals under threat. A joint response from history of science, technology and medicine editors. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):138–141, 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/14/pdf/rhm14-2-edito.pdf.

Proust:2008:QCU

[133] Christine Proust. Quantifier et calculer: usages des nombres à Nippur. (French) [Quantify and calculate: uses of numbers at Nippur]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):143–209 (2009), 2008. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/14/pdf/smf_rhm_14_143-209.pdf.

Alfonsi:2008:EBA

[134] Liliane Alfonsi. Étienne Bézout: analyse algébrique au siècle des Lumières. (French) [Étienne Bézout: algebraic analysis in the century of the Enlightened]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):211–287 (2009), 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/14/pdf/smf_rhm_14_211-287.pdf.

Ehrhardt:2008:EGC

[135] Caroline Ehrhardt. Évariste Galois, un candidat à l’École préparatoire en 1829. (French) [Évariste Galois, a candidate for the Preparatory School in 1829]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 14(2):289–328 (2009), 2008. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/14/pdf/smf_rhm_14_289-328.pdf.

Schappacher:2009:Ea

[136] Norbert Schappacher. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(1):1–3, 2009. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/RevueHistoireMath/15/pdf/rhm15-1-edito.pdf.

Schappacher:2009:Eb

[137] Norbert Schappacher. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(1):5–6, 2009. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/RevueHistoireMath/15/pdf/rhm15-1-edito.pdf.

Audin:2009:PSO

[138] Michèle Audin. Publier sous l’Occupation. I. Autour du cas de Jacques Feldbau et de l’Académie des sciences. (French) [Post under the Occupation. I. Around the case of Jacques Feldbau and the Academy of Sciences]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(1):7–57, 2009. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/15/pdf/smf_rhm_15_7-57.pdf.

Guilbaud:2009:RED

[139] Alexandre Guilbaud and Guillaume Jouve. La résolution des équations aux dérivées partielles dans les Opuscules mathématiques de d’Alembert (1761–1783). (French) [The resolution of the partial differential equations in the Mathematical Functions of d’Alembert (1761–1783)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(1):59–122, 2009. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/15/pdf/smf_rhm_15_59-122.pdf.

Rey:2009:NSM

[140] Javier Docampo Rey. A new source for medieval mathematics in the Iberian Peninsula: the commercial arithmetic in MS 10106 (Biblioteca Nacional, Madrid). Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(1):123–177, 2009. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/15/pdf/smf_rhm_15_123-177.pdf.

With an edited medieval text in Castilian.

Schappacher:2011:Ea

[141] Norbert Schappacher. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(2):179–180, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/15/pdf/rhm15-2-
edito.pdf.

Schappacher:2011:Eb

[142] Norbert Schappacher. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(2):181, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/15/pdf/rhm15-2-
edito.pdf.

Gauthier:2009:GGN

[143] Sébastien Gauthier. La géométrie dans la géométrie des nombres: histoire de discipline ou histoire de pratiques à partir des exemples de Minkowski, Mordell et Davenport. (French) [Geometry in number geometry: history of discipline or history of practice from examples of Minkowski, Mordell and Davenport]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(2):183–230 (2010), 2009. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/15/pdf/smf_rhm_15_183-230.pdf.

Chorlay:2009:PAG

[144] Renaud Chorlay. Passer au global: le cas d’Élie Cartan, 1922–1930. (French) [Going global: the case of Élie Cartan, 1922–1930]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 15(2):231–316 (2010), 2009. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/15/pdf/smf_rhm_15_231-316.pdf.

Gessner:2010:SMI

[145] Samuel Gessner. Savoir manier les instruments: la géométrie dans les écrits italiens d’architecture (1545–1570). (French) [Knowing how to handle instruments: geometry in Italian architectural writings (1545–1570)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(1):1–62, 2010. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/16/pdf/smf_rhm_16_1-62.pdf.

Lubet:2010:CSC

[146] Jean-Pierre Lubet. Calcul symbolique et calcul intégral de Lagrange à Cauchy. (French) [Symbolic calculation and integral calculus of Lagrange to Cauchy]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(1):63–131, 2010. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/16/pdf/smf_rhm_16_63-131.pdf.

Schappacher:2010:Ea

[147] Norbert Schappacher. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(2):133–135, 2010. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/16/pdf/rhm16-2-
edito.pdf.

Schappacher:2010:Eb

[148] Norbert Schappacher. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(2):137–139, 2010. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/16/pdf/rhm16-2-
edito.pdf.

Bullynck:2010:FTN

[149] Maarten Bullynck. Factor tables 1657–1817, with notes on the birth of number theory. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(2):133–216, 2010. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications//RevueHistoireMath/16/pdf/
smf_rhm_16_133-216.pdf.

Voelke:2010:FGS

[150] Jean-Daniel Voelke. Les fondements de la géométrie selon Friedrich Schur. (French) [The foundations of geometry according to Friedrich Schur]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(2):217–286, 2010. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications//
RevueHistoireMath/16/pdf/smf_rhm_16_217-286.pdf.

Volkert:2010:TET

[151] Klaus Volkert. Le tout est-il toujours plus grand que la partie?. (French) [Is everything always bigger than the part?]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 16(2):287–306, 2010. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications//RevueHistoireMath/16/
pdf/smf_rhm_16_287-306.pdf.

Schappache:2011:Ea

[152] Norbert Schappache. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):1–2, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/rhm17-1-
edito.pdf.

Schappache:2011:Eb

[153] Norbert Schappache. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):3–4, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/rhm17-1-
edito.pdf.

Audin:2011:APS

[154] Michèle Audin and Roland Brasseur. Addendum à Publier sous l’occupation I, Revue d’histoire des mathématiques, 15, (2009), 5–57 [MR2582144]. (French) [Addendum to Publish under Occupation I, Revue d’histoire des mathématiques, 15, (2009), 5–57 [MR2582144]]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):5–7, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_5-7.pdf.

Bruneau:2011:LME

[155] Olivier Bruneau. Le De linearum de Maclaurin: entre Newton et Poncelet. (French) [The De linearum of Maclaurin: between Newton and Poncelet]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):9–39, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_9-39.pdf.

Boucard:2011:LPT

[156] Jenny Boucard. Louis Poinsot et la théorie de l’ordre: un chaînon manquant entre Gauss et Galois?. (French) [Louis Poinsot and the theory of order: a missing link between Gauss and Galois?]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):41–138, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_41-138.pdf.

Siegmund-Schultze:2011:OBG

[157] Reinhard Siegmund-Schultze. Opposition to the boycott of German mathematics in the early 1920s: letters by Edmund Landau (1877–1938) and Edwin Bidwell Wilson (1879–1964). Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(1):139–165, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_
17_139-165.pdf.

Schappacher:2011:Ec

[158] Norbert Schappacher. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):167–170, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/rhm17-2-
edito.pdf.

Schappacher:2011:Ed

[159] Norbert Schappacher. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):171–174, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/rhm17-2-
edito.pdf.

Ehrhardt:2011:EGS

[160] Caroline Ehrhardt. Évariste Galois and the social time of mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):175–210, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_175-210.pdf.

Goldstein:2011:CHS

[161] Catherine Goldstein. Charles Hermite’s stroll through the Galois fields. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):211–270, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_211-270.pdf.

Anonymous:2011:CJP

[162] Anonymous. Camille Jordan (1838–1922) painted in 1858 by his uncle Pierre Puvis de Chavanne. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):272, 2011. CO-DEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL

http://smf4.emath.fr/Publications//RevueHistoireMath/17/pdf/
smf_rhm_17_272-272.pdf.

Brechenmacher:2011:SPE

[163] Frédéric Brechenmacher. Self-portraits with Évariste Galois (and the shadow of Camille Jordan). Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):273–371, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/17/pdf/smf_rhm_
17_273-371.pdf.

Archibald:2011:DEA

[164] Tom Archibald. Differential equations and algebraic transcendents: French efforts at the creation of a Galois theory of differential equations 1880–1910. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):373–401, 2011. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/17/pdf/smf_rhm_17_373-401.pdf.

Weber:2011:EGO

[165] Anne-Gaëlle Weber and Andrea Albrecht. Évariste Galois ou le roman du mathématicien. (French) [Évariste Galois or the mathematician’s novel].

Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 17(2):403–435, 2011. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/17/pdf/smf_rhm_17_403-435.pdf.

Anonyme:2012:E

[166] Anonyme. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(1):1–2, 2012. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/18/html/.

Anonymous:2012:E

[167] Anonymous. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(1):3–4, 2012. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/18/html/.

Chambon:2012:NNP

[168] Grégory Chambon. Notations de nombres et pratiques de calcul en Mésopotamie: réflexions sur le système centésimal de position. (French) [Numerical notation and computing practices in M esopotamia: reflections on the centesimal position system]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(1):5–36, 2012. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/18/pdf/smf_rhm_
18_5-36.pdf.

Sarma:2012:ASN

[169] Sreeramula Rajeswara Sarma. The Kaapayādi system of numerical notation and its spread outside Kerala. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(1):37–66, 2012. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/18/pdf/smf_rhm_
18_37-66.pdf.

Herreman:2012:FIG

[170] Alain Herreman. La fonction inaugurale de La géométrie de Descartes. (French) [The inaugural function of Descartes’ Geometry]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(1):67–156, 2012. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/18/pdf/smf_rhm_18_67-156.pdf.

LeMeur:2012:RDQ

[171] Guy Le Meur. Le rôle des diagrammes dans quelques traités de la Petite astronomie. (French) [The role of diagrams in some treatises of the Little Astronomy]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(2):157–221, 2012. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/18/pdf/smf_rhm_18_157-221.pdf.

Clark:2012:PPD

[172] Kathleen M. Clark and Clemency Montelle. Priority, parallel discovery, and pre-eminence: Napier, Bürgi and the early history of the logarithm relation. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(2):223–270, 2012. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/18/pdf/smf_rhm_18_223-270.pdf.

Catellier:2012:EFP

[173] Rémi Catellier and Laurent Mazliak. The emergence of French probabilistic statistics. Borel and the Institut Henri Poincaré around the 1920s. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 18(2):271–335, 2012. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/18/pdf/smf_rhm_18_271-335.pdf.

Ageron:2013:PDS

[174] Pierre Ageron. Le partage des dix-sept chameaux et autres exploits arithmétiques attribués à l’imam Alî. Mouvance et circulation de récits de la tradition musulmane chiite. (French) [The sharing of the seventeen camels and other arithmetic exploits attributed to the Imam Alî. Mouvance and circulation of stories of the Shiite Muslim tradition]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(1):1–41, 2013. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/19/pdf/smf_rhm_19_1-41.pdf.

Ycart:2013:CME

[175] Bernard Ycart. A case of mathematical eponymy: the Vandermonde determinant. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(1):43–77, 2013. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/19/pdf/smf_rhm_19_43-77.pdf.

Verdier:2013:EPV

[176] Norbert Verdier. Éditer puis vendre des mathématiques avec la maison Bachelier (1812–1864). (French) [É diter then sell mathematics with the house Bachelier (1812–1864)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(1):79–145, 2013. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/19/pdf/smf_rhm_
19_79-145.pdf.

Anonyme:2013:D

[177] Anonyme. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(2):147–148, 2013. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/19/pdf/rhm19-2-edito.pdf.

Anonymous:2013:E

[178] Anonymous. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(2):149–150, 2013. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/19/pdf/rhm19-2-edito.pdf.

Herreman:2013:LST

[179] Alain Herreman. L’inauguration des séries trigonométriques dans la Théorie analytique de la chaleur de Fourier et dans la controverse des cordes vibrantes. (French) [The introduction of trigonometric series in Fourier’s Analytical Theory of Heat and in the controversy over vibrating strings]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(2):151–243, 2013. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/19/pdf/smf_rhm_19_151-243.pdf.

Hayashi:2013:GIN

[180] Takao Hayashi. The Gaitapañcaviśī attributed to Śrīdhara. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(2):245–332, 2013. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/19/pdf/smf_rhm_19_245-332.pdf.

Wagner:2013:LAI

[181] Jacques Wagner. Les Lumières et l’astronomie indienne. (French) [The Enlightened and Indian astronomy]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 19(2):333–342, 2013. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/19/pdf/smf_rhm_
19_333-342.pdf.

Anonyme:2014:E

[182] Anonyme. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(1):1–2, 2014. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/20/pdf/rhm20-1-edito.pdf.

Anonymous:2014:E

[183] Anonymous. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(1):3–4, 2014. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/20/pdf/rhm20-1-edito.pdf.

Despeaux:2014:MQC

[184] Sloan Evans Despeaux. Mathematical questions: a convergence of mathematical practices in British journals of the eighteenth and nineteenth centuries. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(1):5–71, 2014. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/20/pdf/smf_rhm_20_5-71.pdf.

Cogliati:2014:RCM

[185] Alberto Cogliati. Riemann’s Commentatio mathematica, a reassessment. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(1):73–94, 2014. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/20/pdf/smf_rhm_20_73-94.pdf.

Eckes:2014:WGP

[186] Christophe Eckes. Weyl et la géométrisation de la physique. (French) [Weyl and the geometrization of physics]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(1):95–170, 2014. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/20/
pdf/smf_rhm_20_95-170.pdf.

Rommevaux-Tani:2014:INI

[187] Sabine Rommevaux-Tani. Irrationalité des nombres, irrationalité des lignes selon Michael Stifel et Simon Stevin. (French) [Irrationality of numbers, irrationality of lines according to Michael Stifel and Simon Stevin].

Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(2):171–209, 2014. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/20/pdf/smf_rhm_20_171-209.pdf.

Lutzen:2014:CAI

[188] Jesper Lützen. 17th century arguments for the impossibility of the indefinite and the definite circle quadrature. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(2):211–251, 2014. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/20/
pdf/smf_rhm_20_211-251.pdf.

Romera-Lebret:2014:NGD

[189] Pauline Romera-Lebret. La nouvelle géométrie du triangle à la fin du XIXe siècle: des revues mathématiques intermédiaires aux ouvrages d’enseignement. (French) [The new triangle geometry at the end of the 19th century: Intermediate mathematical journals to teaching books]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 20(2):253–302, 2014. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/20/pdf/smf_rhm_20_253-302.pdf.

Anonyme:2015:PIB

[190] Anonyme. À propos des indicateurs bibliographiques : la position du comité de rédaction de la Revue d’histoire des mathématiques. (French) [Bibliographic indicators: the Revue d’histoire des mathématiques editorial board’s position]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):1–2, 2015. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/21/pdf/rhm21-1-edito.pdf.

Anonymous:2015:BIR

[191] Anonymous. Bibliographic indicators: the Revue d’histoire des mathématiques editorial board’s position. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):3–4, 2015. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.

emath.fr/RevueHistoireMath/21/pdf/rhm21-1-edito.pdf.

Clain:2015:PAR

[192] Teresa Costa Clain. La Pratica d’arismetica de Ruy Mendes dans le contexte des arithmétiques marchandes ibériques. (French) [The Pratica of Arismetica by Ruy Mendes in the context of Iberian market arithmetics]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):5–39, 2015. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/21/pdf/smf_rhm_21_5-39.pdf.

Roque:2015:LPM

[193] Tatiana Roque. L’originalité Poincaré mécanique céleste: pratique des solutions périodiques dans un réseau de textes. (French) [The Poincaré originality in celestial mechanics: practice of periodic solutions in a text network]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):41–109, 2015. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/21/pdf/smf_rhm_21_41-109.pdf.

Barrow-Green:2015:AAG

[194] June Barrow-Green. “Anti-aircraft guns all day long”: Karl Pearson and computing for the Ministry of Munitions. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):111–150, 2015. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic).

URL http://smf4.emath.fr/Publications/RevueHistoireMath/21/
pdf/smf_rhm_21_111-150.pdf.

Bernard:2015:TBP

[195] Julien Bernard. Les tapuscrits barcelonais sur le problème de l’espace de Weyl. (French) [The typescripts of Barcelona on the problem of Weyl space]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(1):151–171, 2015. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/21/pdf/smf_rhm_21_151-171.pdf.

Petitgirard:2015:LNM

[196] Loïc Petitgirard. L’ingénieur Nicolas Minorsky (1885–1970) et les mathématiques pour l’ingénierie navale, la théorie du contrôle et les oscillations non linéaires. (French) [The engineer Nicolas Minorsky (1885– 1970) and mathematics for naval engineering, control theory and nonlinear oscillations]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(2):173–216, 2015. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/21/pdf/smf_rhm_21_173-216.pdf.

Smadja:2015:SVG

[197] Ivahn Smadja. Sanskrit versus Greek ‘proofs’: history of mathematics at the crossroads of philology and mathematics in nineteenth-century Germany. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 21(2):217–349, 2015. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/21/pdf/smf_rhm_21_217-349.pdf.

Rommevaux-Tani:2016:RPQ

[198] Sabine Rommevaux-Tani. La réception par quelques mathématiciens européens du XVI^{e} siècle des travaux des algébristes italiens sur les équations du troisième degré: réticence de la plupart et avancées significatives de Stevin. (French) [The reception by some European mathematicians of the XVI-th century of the work of Italian algebrists on the equations of the third degree: reticence of most and significant advances of Stevin]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(1):1–52, 2016. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/22/pdf/smf_rhm_22_1-52.pdf.

Eckes:2016:PAC

[199] Christophe Eckes. Un premier aperçu de la correspondance Hecke/Weyl (1930–1938). (French) [A first glimpse of the Hecke/Weyl correspondence (1930–1938)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(1):53–95, 2016. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/22/pdf/smf_rhm_22_53-95.pdf.

Herreman:2016:LCA

[200] Alain Herreman. L’introduction des courbes algébriques par Descartes: genèse et inauguration. Complément à la conjecture de H. Bos sur le rôle historique du problème de Pappus. (French) [The introduction of algebraic curves by Descartes: genesis and inauguration. Complement to the conjecture of H. Bos on the historical role of the Pappus problem]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(1):97–137, 2016. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/22/pdf/smf_rhm_22_97-137.pdf.

Anonyme:2016:E

[201] Anonyme. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(2):??, 2016. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/22/pdf/rhm22-2-edito.pdf.

Anonymous:2016:E

[202] Anonymous. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(2):??, 2016. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/22/pdf/rhm22-2-edito.pdf.

daSilva:2016:CAK

[203] Maria do Céu Peireira da Silva. On the circulation of algebraic knowledge in the Iberian península: the sources of Pérez de Moya’s Tratado de Arithmetica (1573). Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(2):145–184, 2016. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/22/pdf/smf_rhm_22_145-184.pdf.

Preveraud:2016:EUE

[204] Thomas Preveraud. Les États-Unis, espace de compromis: l’adaptation des algèbres françaises de Lacroix et Bourdon aux usages domestiques (1818–1835). (French) [The United States, compromise space: The adaptation of French Lacroix and Bourdon algebras to domestic use (1818– 1835)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(2):185–221, 2016. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/22/pdf/smf_rhm_22_185-221.pdf.

Craik:2016:MAP

[205] Alex D. D. Craik. Mathematical analysis and physical astronomy in great Britain and Ireland, 1790–1831: some new light on the French connection. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 22(2):223–294, 2016. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/22/pdf/smf_rhm_22_223-294.pdf.

Anonyme:2017:E

[206] Anonyme. Éditorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(1):1–2, ???? 2017. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/23/pdf/rhm23-1-edito.pdf.

Anonymous:2017:E

[207] Anonymous. Editorial. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(1):3–4, ???? 2017. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/RevueHistoireMath/23/pdf/rhm23-1-edito.pdf.

Cousin:2017:CNL

[208] Marion Cousin. Sur la création d’une nouvelle langue mathématique japonaise pour l’enseignement de la géométrie élémentaire durant l’ère Meiji (1868–1912). (French) [On the creation of a new Japanese mathematical language for the teaching of elementary geometry during the Meiji era (1868–1912)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(1):5–70, 2017. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.

fr/Publications/RevueHistoireMath/23/pdf/smf_rhm_23_5-70.pdf.

Lorenat:2017:RIE

[209] Jemma Lorenat. Radical, ideal and equal powers: new definitions in modern geometry 1814–1826. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(1):71–123, 2017. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/23/pdf/smf_rhm_23_71-123.pdf.

Haffner:2017:ECC

[210] Emmylou Haffner. Esquisse d’une cartographie des cahiers d’Élie Cartan. (French) [Sketch of a cartography of the Élie Cartan’s notebooks]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(1):125–182, 2017. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/23/pdf/smf_rhm_23_125-182.pdf.

Luciano:2017:CMS

[211] Erika Luciano. Characterizing a mathematical school: Shared knowledge and Peano’s Formulario. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(2):183–231, 2017. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL http://smf4.emath.fr/

Publications/RevueHistoireMath/23/pdf/smf_rhm_23_183-231.pdf.

Moyon:2017:RCO

[212] Marc Moyon. La restauration et la comparaison, ou l’art de résoudre des équations quadratiques dans l’Europe latine. (French) [Restoration and comparison, or the art of solving quadratic equations in the Latin Europe].

Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 23(2):233–299, 2017. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications/
RevueHistoireMath/23/pdf/smf_rhm_23_233-299.pdf.

Le:2018:PEC

[213] François Lê. Le paramétrage elliptique des courbes cubiques par Alfred Clebsch. (French) [Elliptical parameterization of cubic curves by Alfred Clebsch]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 24(1):1–39, 2018. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications//
RevueHistoireMath/24/pdf/smf_rhm_24_1-39.pdf.

Ben-El-Mechaiekh:2018:TDL

[214] Hichem Ben-El-Mechaiekh and Robert W. Dimand. Texts & documents: Louis Bachelier’s 1938 monograph on the calculus of speculation: Mathematical finance and randomness of asset prices in Bachelier’s later work. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 24(1):41–106, 2018. CODEN ???? ISSN 1262-022X (print),

1777-568X (electronic). URL http://smf4.emath.fr/Publications//
RevueHistoireMath/24/pdf/smf_rhm_24_41-106.pdf.

Rommevaux-Tani:2018:TMR

[215] Sabine Rommevaux-Tani. Une théorie de la mesure des rapports dans le Chilias Logarithmorum de Kepler (1624). (French) [A theory of the measurement of relationships in Kepler’s Chilias Logarithmorum (1624)]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 24(2):107–206, ???? 2018. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL https://smf.

emath.fr/publications/une-theorie-de-la-mesure-des-rapports-
dans-le-chilias-logarithmorum-de-kepler-1624.

Morel:2018:FLG

[216] Thomas Morel. Five lives of a Geometria subterranea (1708–1785). Authorship and knowledge circulation in practical mathematics. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 24(2):207–258, ???? 2018. CODEN ???? ISSN 1262-022X (print), 1777-568X (electronic). URL https://smf.emath.fr/

publications/cinq-vies-dune-geometria-subterranea-1708-1785-
autorite-et-circulation-des.

Eckes:2018:ORR

[217] Christophe Eckes. Organiser le recrutement de recenseurs français pour le Zentralblatt à l’automne 1940: les premiers liens entre Harald Geppert, Helmut Hasse et Gaston Julia sous l’Occupation. (French) [Organizing the recruitment of French enumerators for the Zentralblatt in the fall of 1940: the first links between Harald Geppert, Helmut Hasse and Gaston Julia under the Occupation]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 24(2):259–329, ???? 2018. CODEN ???? ISSN 1262-022X (print), 1777-568X (elec-

tronic). URL https://smf.emath.fr/publications/organiser-le-
recrutement-de-recenseurs-francais-pour-le-zentralblatt-lautomne-
1940-les.

Vergnerie:2019:LSF

[218] Cédric Vergnerie. L’algèbre sans les fictions des racines: Kronecker et la théorie des caractéristiques dans les Vorlesungen über die algebraischen Gleichungen. (French) [Algebra without the fictions of the roots: Kronecker and the theory of the characteristics in the Vorlesungen über die algebraischen Gleichungen [Lectures on algebraic equations]]. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 25(1):1–107, ???? 2019. CODEN ???? ISSN 1262-022X (print), 1777-568X

(electronic). URL https://smf.emath.fr/publications/lalgebre-
sans-les-fictions-des-racines-kronecker-et-la-theorie-des-caracteristiques.

Friedman:2019:PNV

[219] Michael Friedman. A plurality of (non)visualizations: Branch points and branch curves at the turn of the 19th century. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 25(1): 109–194, ???? 2019. CODEN ???? ISSN 1262-022X (print), 1777-

568X (electronic). URL https://smf.emath.fr/publications/une-
diversite-de-visualisations-et-non-visualisations-points-de-
branchement-et-courbes.

Boucard:2020:NTN

[220] Jenny Boucard. Number theory in the Nouvelles annales de mathématiques (1842–1927): a case study about mathematical journals for teachers and students. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 26(1):3–72, ???? 2020. ISSN 1262-022X (print), 1777-568X

(electronic). URL https://smf.emath.fr/publications/une-etude-
de-cas-sur-les-journaux-mathematiques-pour-enseignants-et-eleves-
la-theorie.

Freitas:2020:TDC

[221] Pedro J. Freitas. Texts and documents : Charles Hermite’s letters to Francisco Gomes Teixeira. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 26(1):73–95, ???? 2020. ISSN 1262-022X (print), 1777-568X (electronic). URL https://smf.emath.fr/

publications/textes-et-documents-les-lettres-de-charles-hermite-
francisco-gomes-teixeira.

Dickson:1919:HTN

[222] Leonard E. (Leonard Eugene) Dickson. History of the Theory of Numbers, volume 256 (vol. I–III) of Carnegie Institution of Washington publication. Carnegie Institution of Washington, Washington, DC, USA, 1919–1923. ???? pp. LCCN QA241 .D49.

Dickson:1966:HTN

[223] Leonard E. (Leonard Eugene) Dickson. History of the Theory of Numbers. Chelsea Publishing Co., New York, NY, USA, 1966. ???? pp. LCCN QA241 .D5 1966.

Dickson:2005:HTN

[224] Leonard E. (Leonard Eugene) Dickson. History of the Theory of Numbers. Dover books on mathematics. Dover Publications, Inc., New York, NY, USA, 2005. ISBN 0-486-44232-2 (vol. 1), 0-486-44233-0 (vol. 2), 0-486-44234-9 (vol. 3). xvi + 486 (vol. 1); xxii + 803 (vol. 2); v + 313 (vol.

3) pp. LCCN QA241 .D5 2005. URL http://www.loc.gov/catdir/
enhancements/fy0623/2005041272-d.html;   http://www.loc.gov/
catdir/enhancements/fy1318/2005041272-t.html.